A box contains eight chocolates -- four with pecans, three with coconut, and one with raspberry creme. Harley Quinn eats three of the chocolates, chosen randomly. If it is known that at least one of the ones she eats has coconut, what's the probability that exactly one of them has coconut?

Enter your answer as a fraction in lowest terms

Answer by Alex.33(110)

Now we know that at least one of the ones Ms.Quinn eats has coconut.
number of possible cases:
1 coconut:3c1*5c2/=30
2 coconut:3c2*5c1=15
3 cocolut:3c3=1.
overall: 30+15+1=46.
P(1 chocolate)=30/46=15/23.
Hope this solves your problem. This is the basic approach towards such problems.
Tip: When doing combinations, BE VERY CAREfUL NOT TO DOUBLE COUNT ANY CASE.

If you're stuck(like with 2 methods both of which seem right but to different results), a quick way out is to list all the cases possible. This can 100% help you spot your problem and avoid it next time.